OpenStudy (anonymous):

Which of the following equations have the same solution? Solve all three equations separately, and explain each step.

OpenStudy (anonymous):

I. 4x – 14 = 2x + 8 II. 3x – 21 = 12 III. the quantity of x minus 2, divided by 5 equals the quantity of 4x plus 2, divided by 25

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Equation I = 11 Equation II = 11 Equation III = 12

OpenStudy (anonymous):

steps?

OpenStudy (anonymous):

u there?

OpenStudy (igreen):

Here's my work: \(4x – 14 = 2x + 8\) \(4x = 2x + 22\) \(2x = 22\) \(x = 11\) \(3x – 21 = 12\) \(3x = 33\) \(3x = 11\) \(\Large\frac{x-2}{5} = \frac{4x+2}{25}\) Cross Multiply: \(25x - 50 = 20x + 10\) \(25x = 20x + 60\) \(5x = 60\) \(x = 12\)